OpenStudy (anonymous):

Find another solution if y1=cosh(x) is a solution to y''-y=0 using reduction of order method

OpenStudy (anonymous):

let y=u(x)y1 y=ucosh(x) y'=usinh(x)+u'cosh(x) y''=ucosh(x)+u'sinh(x)+u'sinh(x)+u''cosh(x) y''-y=0=ucosh(x)+2u'sinh(x)+u''cosh(x)-ucosh(x)=2u'sinh(x)+u''cosh(x)

OpenStudy (anonymous):

letting w=u' , you get the linear equation w'cosh(x)-2sinh(x)w=0

jimthompson5910 (jim_thompson5910):

Did you manage to find u(x) or w(x)?

OpenStudy (anonymous):

i feel stupid -.- i solved to get u but never got y

OpenStudy (anonymous):

i ended up with \[cosh^2(x)w=c\] =\[c\int{}{}sec^2h(x)\]=ctanh(x)

jimthompson5910 (jim_thompson5910):

don't be, I'm more lost than you are most of the time lol

jimthompson5910 (jim_thompson5910):

so w(x) = tanh(x)?

OpenStudy (anonymous):

y=ucosh(x)=tanh(x)cosh(x)=sinh(x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the constant isn't really need since you're just trying to find the general

jimthompson5910 (jim_thompson5910):

looks good and it looks like you definitely know what you're doing

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

no i didn't i totally was sitting here for hours trying to figure out what i did wrong but i hadn't done anything wrong. I just forgot to solve for y. I had u

jimthompson5910 (jim_thompson5910):

lol glad you found the culprit

OpenStudy (anonymous):

onto chapter 4.3 lol

OpenStudy (mendicant_bias):

I don't understand how, in your second post, one of the terms became negative. How did 2sinh(x)mu'(x) become negative after w(x) was substituted for mu'(x)?